OpenStudy (annjewell1998):

Solve the equation for y: 4x + 3y = -12.

OpenStudy (anonymous):

Would you like some help?

OpenStudy (annjewell1998):

Yes please.

OpenStudy (anonymous):

Okay, do you know what we are subtracting?

OpenStudy (annjewell1998):

3y?

OpenStudy (anonymous):

No. Not quite. We are solving for y. Therefore, we subtract 4x from both sides.

OpenStudy (annjewell1998):

Can I get back on later, i have to go to my next class, i'll be on during my lunch though.

OpenStudy (anonymous):

No problem. Pm me when you need me. :)

OpenStudy (anonymous):

you need help

OpenStudy (anonymous):

Should i write the steps how to solve it

OpenStudy (calculusxy):

\[4x + 3y = -12 \rightarrow 3y = -12 - 4x \rightarrow \frac{ 3y }{ 3 } = \frac{ -12 - 4x }{ 3 } \rightarrow y = -4 - \frac{ 4x }{ 3 }\]

OpenStudy (calculusxy):

That's my work, I am not sure if it's correct :\

OpenStudy (anonymous):

right

OpenStudy (calculusxy):

@Abdullahbasra007 I am right?

OpenStudy (annjewell1998):

Well the answer choices were a y = -12 - 4/3 x b y = -4 - 4/3 x c y = -12 + 4x d y = -4 + 4x

OpenStudy (annjewell1998):

4x + 3y = -12 -4x -4x 3y= -12-4x this is as far as i get and i don't know where to go next.

OpenStudy (calculusxy):

So the answer that I had can also be written as: \[\color \green{y = -4 - \frac{ 4 }{ 3 }x}\]

OpenStudy (annjewell1998):

Thank you, but how did you get that? can you explain the steps please?

OpenStudy (calculusxy):

Well what I did was isolate y. That's why I moved 4x to the other side and then divided that right side all by 3.

OpenStudy (annjewell1998):

Oh okay, thank you.